# Self-Admitted GenAI Usage in Open-Source Software

### v2

@@ Abstract
The widespread adoption of generative AI (GenAI) tools such as GitHub Copilot and ChatGPT is transforming software development. Since generated source code is virtually impossible to distinguish from manually written code, their real-world usage and impact on open-source software development remain poorly understood. In this paper, we introduce the concept of self-admitted GenAI usage, that is, developers explicitly referring to the use of GenAI tools for content creation in software artifacts. Using this concept as a lens to study how GenAI tools are integrated into open-source software projects, we analyze a curated sample of more than 250,000 GitHub repositories, identifying 1,292 such self-admissions across 156 repositories in commit messages, code comments, and project documentation. Using a mixed methods approach, we derive a taxonomy of 32 tasks, 10 content types, and 11 purposes associated with GenAI usage based on 284 qualitatively coded mentions. We then analyze 13 documents with policies and usage guidelines for GenAI tools and conduct a developer survey to uncover the ethical, legal, and practical concerns behind them. Our findings reveal that developers actively manage how GenAI is used in their projects, highlighting the need for project-level transparency, attribution, and quality control practices in the new era of AI-assisted software development. Finally, we examine the impact of GenAI adoption on code churn in 151 repositories with self-admitted GenAI usage and find no general increase, contradicting popular narratives on the impact of GenAI on software development.
Index Terms:
Software Engineering, Generative Artificial Intelligence, Large Language Models, Software Maintenance and Evolution
@@ 1 Introduction
Understanding how developers integrate GenAI tools into their workflows is essential to assess their practical impact on software development processes. Open-source software projects, with their collaborative nature and publicly accessible repositories, offer a unique context for studying this integration [9]. Developers often document their activities through commit messages, code comments, and project documentation, potentially providing valuable insights into how GenAI tools are perceived, employed, and acknowledged in real-world scenarios.
@@
We conducted a large-scale empirical study of over 250,000 open-source software repositories hosted on GitHub. Our investigation focused on identifying explicit mentions of GenAI tools in various project artifacts and analyzing how these mentions relate to development activities. We followed a mixed methods approach, combining a qualitative analysis of GenAI-related mentions with a quantitative examination of the code churn over time.
This paper makes three key contributions:
@@
    We present a taxonomy of 32 development tasks, 10 content types, and 11 usage purposes, derived from a qualitative analysis of 1,292 mentions of GenAI usage in GitHub repositories.
@@
    We provide actionable recommendations for responsible and transparent GenAI tool usage in open-source projects based on an analysis of 13 GenAI-related policies and recommendations, and a developer survey.
@@
    We reveal diverse patterns of how GenAI adoption impacts code churn, challenging claims regarding the impact of GenAI, based on a analysis of code churn pre- and post-GenAI adoption in 151 GitHub repositories.
@@ 2 Methodology
We followed a mixed-methods research design. After retrieving instances of self-admitted GenAI usage from open-source GitHub repositories, we conducted a qualitative analysis to answer RQ1. Through multiple iterative coding phases, we labeled these instances to classify supported tasks and generated content. Since this qualitative analysis yielded a considerable number of statements that focused on the regulation or recommendation of GenAI practices, we followed up with a closer analysis of these aspects as part of RQ2. For RQ3, we used self-admitted GenAI usages to approximate the time when the projects started using GenAI tools, to analyze the effect of GenAI usage on code churn using a Regression Discontinuity Design (RDD).
@@ 2.1 Repository Sampling
The foundation of our research is a large sample of open-source GitHub repositories. We first selected 735,669 repositories using the GitHub search tool provided by Dabic et al. [12]. We selected repositories primarily written in the five most popular programming languages identified in a 2024 GitHub report [13]: Python, JavaScript, TypeScript, Java, and C#. Since RQ3 aims at a comparison of code churn before and after projects started using GenAI tools, we only selected repositories that: (1) were created before the ChatGPT launch date (30 November 2022) and (2) had at least one commit on or after this date. Moreover, to eliminate duplicates, we excluded forks. Our initial sample of GitHub projects contained 258,216 repositories distributed across Python (77,542), JavaScript (48,500), TypeScript (37,424), Java (25,160), and C# (18,436).
@@
To select engineered software projects with sufficient development data, we excluded repositories in the first quartile ($Q_{1}$) for each metric, therefore removing the lowest 25%. Furthermore, we excluded repositories with a code ratio (defined as $lines\_of\_code/(lines\_of\_code+lines\_of\_comments)$) outside the 97% confidence interval. The rationale behind this threshold is that engineered software projects are usually documented using source code comments. Filtering out repositories beyond the 97% confidence interval helps eliminate outliers, that is, repositories with very little code, or codebases dominated by code without comments. A manual review further confirmed that this ratio serves as a reliable indicator for filtering out non-software or poorly structured projects.
@@ 2.2 Identifying Self-Admitted GenAI Usages
To identify self-admitted GenAI usage in our filtered sample of GitHub repositories, we retrieved mentions of the two most popular GenAI tools among developers [15]: ChatGPT and GitHub Copilot. Then, in the second step, we annotated these mentions to identify those related to content generation. We wrote a Python script for the following process:
@@
where the variable llm_tool was assigned the value r’chat[ \-_]{0,1}gpt’ for ChatGPT and r’co[ \-_]{0,1}pilot’ for GitHub Copilot. These patterns allowed us to capture variations in how these tools were referenced while minimizing false positives. We developed heuristics to further reduce the number of false positives. For example, we noticed that in false positive matches, the mentions of GenAI tools were often surrounded by commas or underscores, for example, when they were part of URLs for API calls. Our supplementary material contains the full source code that documents our retrieval approach.
Running the above retrieval process on all repos yielded 3,004 mentions of GenAI tools: 1,572 in commit messages, 397 in source code comments, and 1,035 in documentation files. These mentions were automatically obtained using regular expressions and filtered according to heuristics. However, they still included mentions that were not related to content generation. Thus, we conducted a thorough manual inspection of all mentions to eliminate false positives. This review process was guided by the following instructions:
@@
To evaluate the coding instructions, two authors independently labeled a sample of mentions, deciding whether they should be included or not. We calculated a sample size of 341 mentions (of 3,004) to achieve estimates with a 95% confidence level and a 5% confidence interval. The inspection resulted in disagreement between the two authors for only 14 cases (4% of the sample). The two authors discussed these cases and tried to reach a consensus. During these discussions, a third author helped resolve each disagreement and suggested possible improvements to the categories. To assess inter-rater reliability, we computed Fleiss’ kappa [16] by applying bootstrap resampling methods with 1,000 iterations. The resulting 95% confidence interval was estimated to be (0.87, 0.95), indicating an almost perfect agreement. Given this high agreement, the first author continued to inspect the remaining mentions alone. In total, we identified 1,292 true-positive mentions of GenAI tools that were aligned with our inclusion criteria. We found true-positive mentions in 156 repositories (11 Python, 12 JavaScript, 37 TypeScript, 49 Java, 47 TypeScript repositories).
@@ 3 Reasons for Mentioning GenAI Tools
To answer RQ1, we qualitative analyzed the GenAI mentions that we collected and curated, categorizing them according to tasks, contents, and purposes.
@@ 3.1 Method
A preliminary analysis revealed that 1,008 mentions were from Copilot-generated commit messages created in context of pull requests in a single repository named pancakeswap/pancake-frontend. Given this overrepresentation of one repository and GenAI mention type, we set these mentions aside during the initial round of coding to avoid skewing the development of the coding schema. After establishing a stable set of categories through analysis of the remaining mentions, we returned to initially deferred cases for subsequent review and integration.
To build the code book, two authors independently analyzed 284 GenAI mentions. The categorization and code book development was guided by the following questions:
@@
Using Fleiss’ kappa [16], we assessed the interrater reliability between the two coders. The analysis yielded “substantial” to “almost perfect” agreement levels on task ($k=0.81-0.89$), content ($k=0.95-0.99$), and purpose ($k=0.79-0.92$), according to standard guidelines for interpreting $k$ [19]. Through iterative discussions, the two coders worked to achieve consensus on the categorizations, with a third researcher arbitrating unresolved disagreements and recommending refinements to the categories. Based on this strong level of agreement, the first author independently coded the 1,008 mentions that we had initially deferred.
@@ 3.2 Results
Table II: GenAI-assisted tasks (RQ1): Definition and frequency of categories and codes.
@@
Generation PR description Using GenAI tools to create detailed and clear descriptions for pull requests, outlining changes made and their impact to assist reviewers in understanding the modifications. 1,009
Code Using GenAI tools to understand programming tasks described in natural language and generate syntactically correct and logically coherent code snippets. 105
@@
Test data Using GenAI tools to automatically create test input and output based on a given set of software requirements or existing codebase. 9
@@
Comment Using GenAI tools to generate explanatory code comments, enhancing readability and maintainability by clarifying the purpose and logic of code blocks. 9
@@
Test file Using GenAI tools to automatically create test cases and scripts based on a given set of software requirements or existing codebase. 8
@@
Regex Using GenAI tools to craft regular expressions tailored to specific text matching needs, simplifying data validation, extraction, or search tasks. 6
README Using GenAI tools to create initial README documents for projects, providing essential information such as project descriptions, installation instructions, and usage examples. 4
@@
Dummy text Using GenAI tools to produce placeholder text that mimics real content in style, structure, and format. 4
@@
Test method Using GenAI tools to automatically create test cases and scripts based on a given set of software requirements or existing codebase. 2
Code review Using GenAI tools to examine code, suggest improvements, and identify potential issues such as bugs, inefficiencies, or deviations from best practices. 2
Commit
message
Using GenAI tools to generate concise and informative commit messages that summarize code changes, facilitating better version control and project tracking. 2
@@
Tutorial Using GenAI tools to produce instructional content on specific topics, providing step-by-step guidance to help newcomers understand the project. 2
@@
Zod schema Using GenAI tools to generate Zod schemas for TypeScript and JavaScript to enforce type safety and data validation. 2
@@
Test class Using GenAI tools to automatically create test classes based on a given set of software requirements or existing codebase. 1
Coding
practices
Using GenAI tools to generate guidelines and best practices for coding to promote code quality, maintainability, and adherence to industry standards. 1
@@
Variable Using GenAI tools to suggest meaningful and contextually appropriate variable names, improving code semantics and readability. 1
@@
Changelog Using GenAI tools to compile and format changelogs that document changes, features, and fixes in new software versions, enhancing transparency and user communication. 1
@@
Configuration Using GenAI tools to generate project configuration files (e.g., to optimize performance and security settings, user preferences). 1
Text Using GenAI tools to generate general text that is not mentioned above. 8
Translation Text Using GenAI tools to cover text from one language to another, aiming to preserve the original meaning. This task includes the software internationalization (i18n). 49
@@
Code Using GenAI tools to translate code from one programming language to another, maintaining the original logic and functionality while adapting to the syntax and idiomatic patterns of the target language. 1
Optimization Code
refactoring
@@
Using GenAI tools to restructure existing code without altering its functionality, aiming to make the code easier to maintain and extend. 29
Code
improvement
Using GenAI tools to improve existing code, mention is accompanied by “improve”. 5
Maintenance Label
revision
@@
Using GenAI tools to analyze, update, and improve text labels, ensuring clarity, accuracy, and consistency. 8
README
revision
@@
Using GenAI tools to analyze, update, and improve README files, ensuring clarity, accuracy, and consistency. 7
Documentation
revision
@@
Using GenAI tools to analyze, update, and improve technical documents, ensuring clarity, accuracy, and consistency. 4
Changelog
revision
@@
Using GenAI tools to analyze, update, and improve changelogs, ensuring clarity, accuracy, and consistency. 2
Prompt
refinement
Using GenAI tools to optimize and clarify the prompts in the project, ensuring that they are precise, contextually appropriate, and designed to elicit the most relevant and accurate responses. 1
Color
suggestion
@@
Using GenAI tools to suggest color schemes for UI/UX design based on best practices, user preferences, or specific design requirements. 1
@@
Using GenAI tools to analyze software dependencies and suggest updates to ensure compatibility and security while minimizing breaking changes. 1
Version
update
Using GenAI tools to manage and suggest appropriate version numbering for software releases, ensuring systematic and meaningful version control. 1
Comment
revision
Using GenAI tools to analyze, update, and improve source code comments. 1
Other - Using GenAI tools to operate general functionality, like Q&A, blog generation, or unspecific tasks. 12

### v3

@@ Abstract
The widespread adoption of generative AI (GenAI) tools such as GitHub Copilot and ChatGPT is transforming software development. Since generated source code is virtually impossible to distinguish from manually written code, their real-world usage and impact on open-source software development remain poorly understood. In this paper, we introduce the concept of self-admitted GenAI usage, that is, developers explicitly referring to the use of GenAI tools for content creation in software artifacts. Using this concept as a lens to study how GenAI tools are integrated into open-source software projects, we analyze a curated sample of more than 250,000 GitHub repositories, identifying 1,292 such self-admissions across 156 repositories in commit messages, code comments, and project documentation. Using a mixed methods approach, we derive a taxonomy of 32 tasks, 10 content types, and 11 purposes associated with GenAI usage based on 1,292 qualitatively coded mentions. We then analyze 13 documents with policies and usage guidelines for GenAI tools and conduct a developer survey to uncover the ethical, legal, and practical concerns behind them. Our findings reveal that developers actively manage how GenAI is used in their projects, highlighting the need for project-level transparency, attribution, and quality control practices in the new era of AI-assisted software development. Finally, we examine the longitudinal impact of GenAI adoption on code churn in 151 repositories with self-admitted GenAI usage and find no general increase, contradicting popular narratives on the impact of GenAI on software development.
@@ 1 Introduction
As a result, much of what we know about GenAI usage in software development is inferred indirectly, either from vendor-controlled telemetry, laboratory studies, or analyses of repository activity that rely on aggregated metrics rather than direct evidence of GenAI use. This makes it difficult to understand how GenAI tools are actually integrated into collaborative development workflows, how their use is governed in practice, and how their impact should be interpreted at the project level. In real-world settings, developers must decide how much to rely on or revise AI-generated content, and maintainers must determine whether to prohibit, restrict, or encourage GenAI use. Researchers increasingly rely on aggregate metrics such as code churn to assess claims about software quality degradation. Without observable, project-level signals of GenAI usage embedded in software artifacts, these decisions risk being shaped by assumptions, anecdotal evidence, or broad industry narratives rather than by empirical data. Studying explicit references to GenAI usage offers a way to ground these discussions in real development practices, making it possible to examine not only what GenAI is used for, but also how it is acknowledged, regulated, and followed by human action in open-source software projects.
Open-source software (OSS) projects, with their collaborative nature and publicly accessible repositories [6], offer a unique context to study the adoption of GenAI tools. Although such tools promise to support OSS projects by automating development tasks, there are also reports of “AI slop” wasting valuable time of maintainers [41] or GenAI-generated contributions leading to more rework [10].
@@
We conducted a large-scale empirical study of more than 250,000 OSS repositories hosted on GitHub. Our investigation focused on identifying explicit mentions of GenAI tools in various artifacts and analyzing how these mentions relate to development activities. We followed a mixed methods approach, combining a qualitative analysis of GenAI-related mentions with a quantitative examination of code churn over time, resulting in four main contributions:
@@
    We introduce self-admitted GenAI usage as an empirical lens for studying GenAI adoption in open-source software and curate a dataset of 1,292 self-admitted GenAI usages across 156 GitHub repositories.
@@
    Using a mixed-methods approach, we derive a taxonomy of GenAI usage comprising 32 development tasks, 10 content types, and 11 purposes, grounded in a qualitative analysis of the identified usages.
@@
    We empirically analyze how open-source projects govern GenAI usage by examining 13 policies and guidelines, contextualizing them through a developer survey.
  4. 4.
    We assess the longitudinal impact of GenAI adoption on software evolution using a repository-level analysis of code churn in 151 projects, showing that GenAI adoption does not lead to a general increase in churn and that effects are stronger for generation tasks.
@@ 2 Methodology
Figure 1: Overview of the data collection process used to answer our three research questions, from the selection (1) and filtering (2) of GitHub repositories to the extraction of GenAI mentions (3) and the identification of self-admitted GenAI usage (4) in these repositories.
We followed a mixed-methods research design. Our data collection process is visualized in Figure 1. After retrieving instances of self-admitted GenAI usage from open-source GitHub repositories, we conducted a qualitative analysis to answer RQ1. Through multiple iterative coding phases, we labeled these instances to classify supported tasks and generated content. Since this qualitative analysis yielded a considerable number of statements that focused on the regulation or recommendation of GenAI practices, we followed up with a closer analysis of these aspects as part of RQ2. For RQ3, we used self-admitted GenAI usages to approximate the time when the projects started using GenAI tools, to analyze the effect of GenAI usage on code churn using a Regression Discontinuity Design (RDD).
@@ 2.1 Repository Sampling
The foundation of our research is a large sample of open-source GitHub repositories. We selected GitHub as our study platform because it is the largest and most widely used open-source hosting service, with over 500 million repositories according to the 2024 Octoverse report [12], making it the most suitable environment for analyzing trends of GenAI usage in open-source software development. Using the GitHub search tool provided by Dabic et al. [7], we selected repositories primarily written in the five most popular programming languages as of the above-mentioned report [12]: Python, JavaScript, TypeScript, Java, and C#. This focus on the most popular languages ensures that our study is both manageable and relevant to the most commonly used development ecosystems. Since RQ3 aims at a comparison of code churn before and after projects started using GenAI tools, we only selected repositories that: (1) were created before the ChatGPT launch date (30 November 2022) and (2) had at least one commit on or after this date. Moreover, to eliminate duplicates, we excluded forks. Our initial sample of GitHub projects contained 258,216 repositories distributed across Python (77,542), JavaScript (48,500), TypeScript (37,424), Java (25,160), and C# (18,436).
@@
To select engineered software projects with sufficient development data, we excluded repositories in the first quartile ($Q_{1}$) for each metric, therefore removing the lowest 25%. Furthermore, we excluded repositories with a code ratio (defined as lines_of_code/(lines_of_code+lines_of_comments)lines\_of\_code/(lines\_of\_code+lines\_of\_comments)) outside the 97% confidence interval. The rationale behind this threshold is that engineered software projects are usually documented using source code comments. Filtering out repositories beyond the 97% confidence interval helps eliminate outliers, that is, repositories with very little code, or codebases dominated by code without comments. A manual review further confirmed that this ratio serves as a reliable indicator for filtering out non-software or poorly structured projects.
@@ 2.2 Identifying Self-Admitted GenAI Usages
To identify self-admitted GenAI usage in our filtered sample of GitHub repositories, we retrieved mentions of the two most popular GenAI tools among developers as of the 2023 Stack Overflow Developer Survey [40]. In that survey, ChatGPT was identified as the most popular general AI tool and GitHub Copilot as the most popular AI developer tool. Then, in the second step, we annotated these mentions to identify those related to content generation. We wrote a Python script for the following process:
@@
where the variable llm_tool was assigned the value r’chat[ \-_]{0,1}gpt’ for ChatGPT and r’co[ \-_]{0,1}pilot’ for Copilot. These patterns allowed us to capture variations in how these tools were referenced while minimizing false positives. We developed heuristics to further reduce the number of false positives. For example, we noticed that in false positive matches, the mentions of GenAI tools were often surrounded by commas or underscores, for example, when they were part of URLs for API calls. Our supplementary material contains the full source code that documents our retrieval approach.
Running the above retrieval process on all repositories yielded 3,004 mentions of GenAI tools: 1,572 in commit messages, 397 in source code comments, and 1,035 in documentation files. These mentions were automatically obtained using regular expressions and filtered according to heuristics. However, they still included mentions that were not related to content generation. Thus, we conducted a thorough manual inspection of all mentions to eliminate false positives. This review process was guided by the following instructions:
@@
To evaluate the coding instructions, two authors independently labeled a sample of mentions, deciding whether they should be included or not. We calculated a sample size of 341 mentions (of 3,004) to achieve estimates with a 95% confidence level and a 5% confidence interval. The inspection resulted in disagreement between the two authors for only 14 cases (4% of the sample). The two authors discussed these cases and tried to reach a consensus. During these discussions, a third author helped resolve each disagreement and suggested possible improvements to the categories. To assess inter-rater reliability, we computed Fleiss’ kappa [9] by applying bootstrap resampling methods with 1,000 iterations. The resulting 95% confidence interval was estimated to be (0.87, 0.95), indicating an almost perfect agreement. Given this high agreement, the first author continued to inspect the remaining mentions alone. In total, we identified 1,292 true-positive mentions of GenAI tools that were aligned with our inclusion criteria. We found true-positive mentions in 156 repositories (11 Python, 12 JavaScript, 37 TypeScript, 47 C#, and 49 Java repositories).
@@ 3 Reasons for Mentioning GenAI Tools
To answer RQ1, we qualitatively analyzed the GenAI mentions that we collected and curated, categorizing them according to tasks, contents, and purposes.
@@ 3.1 Method
A preliminary analysis revealed that 1,008 mentions were from Copilot-generated commit messages created in the context of pull requests in a single repository named pancakeswap/pancake-frontend. Given this overrepresentation of one repository and GenAI mention type, we set these mentions aside during the initial round of coding to avoid skewing the development of the coding schema. After establishing a stable set of categories through analysis of the remaining mentions, we returned to initially deferred cases for subsequent review and integration.
To build the code book, two authors independently analyzed 284 GenAI mentions. The categorization and code book development were guided by the following questions:
@@
Using Fleiss’ kappa [9], we assessed the interrater reliability between the two coders. The analysis yielded “substantial” to “almost perfect” agreement levels on task (k=0.810.89k=0.81-0.89), content (k=0.950.99k=0.95-0.99), and purpose (k=0.790.92k=0.79-0.92), according to standard guidelines for interpreting $k$ [45]. Through iterative discussions, the two coders worked to achieve consensus on the categorizations, with a third researcher arbitrating unresolved disagreements and recommending refinements to the categories. The first author then independently checked the 1,008 mentions that we had initially deferred.
@@ 3.2 Results
Table II: GenAI-assisted tasks (RQ1): Definition and frequency of categories and codes (n = 284 + 1,008 = 1,292); the code PR description is counted and discussed separately because it only occurred in one repository (see Section 3.2.1).
@@
Generation Code Understand coding tasks written in natural language and generate corresponding code. 105
@@
Test data Create test input/output based on software requirements or the existing codebase. 9
@@
Comment Generate code comments that explain the purpose and logic of code blocks. 9
@@
Test file Create test cases and scripts based on software requirements or the existing codebase. 8
@@
Regex Craft regular expressions tailored to specific text matching needs. 6
README Create README files that provide essential information, e.g., project descriptions. 4
@@
Dummy text Produce placeholder text that mimics real content in style, structure, and format. 4
@@
Test method Create test cases and scripts based on software requirements or the existing codebase. 2
Code review Generate reviews that suggest improvements and identify potential issues in code changes. 2
Commit message Generate commit messages that summarize code changes. 2
@@
Tutorial Produce instructional content on specific topics and step-by-step guidance for projects. 2
@@
Zod schema Create Zod schemas in TypeScript and JavaScript for type safety and data validation. 2
@@
Test class Create test classes based on software requirements or the existing codebase. 1
Coding practices Generate guidelines and best practices for coding in the projects. 1
@@
Variable Suggest meaningful variable names that improve code semantics and readability. 1
@@
Changelog Compile changelogs that document changes, features, and fixes in new software versions. 1
@@
Configuration Generate project-specific configuration files, e.g., performance and security settings. 1
Text Generate general text that is not mentioned above. 8
PR description Create explicit PR descriptions to assist understanding the changes in the PRs. 1,008
Translation Text Convert text between different languages, e.g., software internationalization (i18n). 49
@@
Code Convert code between different programming languages, preserving the original logic and functionality while adapting to the syntax and idiomatic patterns of the target language. 1
@@
Optimization Code refactoring Restructure code without altering its functionality, aiming to make the code maintainable. 29
Code improvement Improve existing code, mention is accompanied by “improve.” 5
@@
Maintenance Label revision Analyze, update, and improve text labels, ensuring clarity, accuracy, and consistency. 8
@@
README revision Analyze, update, and improve README files, ensuring clarity, accuracy, and consistency. 7
@@
Document revision Analyze, update, and improve documents, ensuring clarity, accuracy, and consistency. 4
@@
Changelog revision Analyze, update, and improve changelogs, ensuring clarity, accuracy, and consistency. 2
Prompt refinement Optimize and clarify the prompts to elicit the most relevant and accurate responses. 1
@@
Color suggestion Suggest color schemes for UI/UX design based on best practices and design requirements. 1
@@
Analyze software dependencies and suggest updates to ensure compatibility and security while minimizing breaking changes. 1
Version update Suggest meaningful version numbering for software releases for systematic version control. 1
Comment revision Analyze, update, and improve source code comments. 1
Other - Operate general functionality, like Q&A, blog generation, or unspecific tasks. 12